OpenStudy (anonymous):

Factor Completely: 25x2 − 9

OpenStudy (anonymous):

(5x+3)(5x-3)=0

OpenStudy (anonymous):

Use difference of 2 squares \(a^2 - b^2 = (a+b)(a-b)\) Now, put a = 5x and b=3 into the identity 25x^2 − 9 = (5x-3)(5x+3)

OpenStudy (anonymous):

Thank You (:

OpenStudy (anonymous):

@some1 Please be aware that it's NOT an equation